The number of radial nodes in 3s and 2p respectively are
a. 2 and 0
b. 0 and 2
c. 1 and 2
d. 2 and 1

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
The number of radial nodes is given by expression (n – l - 1). For 3s, number of nodes = 3 – 0 – 1 = 2 For 2p number of nodes = 2 – 1 – 1 = 0